OpenStudy (gabyramirezzz):

Which function's graph has a period of 2? a) y= -4 sin2x b) y= 3 cosx c) y= 2 sinpix d) y= cos(x-pi/2)

OpenStudy (zzr0ck3r):

a) (2pi)/2 = pi b) 2pi/1 = 2pi c) ?

OpenStudy (gabyramirezzz):

2 sine*pi*x

OpenStudy (zzr0ck3r):

\(\sin(ax+c)\) then period is given by \(\frac{2\pi}{a}\)

OpenStudy (zzr0ck3r):

what is the period for c) ?

OpenStudy (gabyramirezzz):

i dont know

OpenStudy (zzr0ck3r):

I showed you how.

OpenStudy (gabyramirezzz):

would it be 2pi/2?

OpenStudy (zzr0ck3r):

=correct

OpenStudy (zzr0ck3r):

no

OpenStudy (zzr0ck3r):

\(\sin(\pi x)=\sin(\pi x+0)\) \(\sin(ax+b)\) What number on the first equation corresponds with \(a\) in the second equation?

OpenStudy (gabyramirezzz):

-4?

OpenStudy (zzr0ck3r):

\(\sin(\pi x+0)\\\sin(ax+b)\) what is \(a\) in the second one?

OpenStudy (gabyramirezzz):

pi

OpenStudy (zzr0ck3r):

period = \(\frac{2\pi}{a}\)

OpenStudy (zzr0ck3r):

what is period?

OpenStudy (gabyramirezzz):

2

OpenStudy (zzr0ck3r):

correct

OpenStudy (gabyramirezzz):

so its a?

OpenStudy (gabyramirezzz):

no

OpenStudy (gabyramirezzz):

wait its c

OpenStudy (zzr0ck3r):

correct

OpenStudy (gabyramirezzz):

yay! thank you

OpenStudy (zzr0ck3r):

np